OpenStudy (anonymous):

A theater has 33 seats in the first row, 35 seats in the second row, 37 seats in the third row, and so on. a. Can the number of seats in each row be modeled by an arithmetic or geometric sequence b. write the general term for a sequence an that gives the number of seats in row n. c. how many seats are there in row 17

OpenStudy (anonymous):

@satellite73 would it be an=33+(n-1)*2 row 17 has 65 chairs?

OpenStudy (anonymous):

or @Directrix is that right, I remember doing this with you.- I believe.

OpenStudy (anonymous):

Mind if I give it a crack? The number it increasing by each row is 2 right?

OpenStudy (anonymous):

yeah so it would be 33+(17-1)*2

OpenStudy (anonymous):

ultimately giving us 17=33+32 17=65 right?

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

holy crap! I was right! LOL

Directrix (directrix):

@PECKCE I remember this. I think we were looking for the number of seats in row 19 on that other problem. >>is that right, I remember doing this with you.- I believe.

OpenStudy (anonymous):

yeah it was and we figured out- I looked back to what we did for finding row 19! you are AWESOME!